News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 16 March 2012 9.00am EDT

During the last week the particle radiation due to Solar activity following the Mar 7 safing event subsided enough to resume observing. Real-time procedures were executed on Mar 12 to verify that the radiation environment as observed by the HRC anticoincidence shield had decreased to an acceptable level and to start a collection of ACIS CCD charge-transfer inefficiency data. Loads for a replanned schedule were uplinked on Mar 13 with 440.0ks of science loss. The observing schedule was interrupted again by a high-radiation event that activated the science instrument safing sequence (SCS 107) and stopped observing at 6:41pm EDT on Mar 13. All spacecraft actions were nominal. A real-time procedure was executed on Mar 14 to verify that the radiation environment as observed by the HRC anticoincidence shield had decreased to an acceptable level. Loads for a replanned schedule were uplinked on Mar 14 with 53.3ks of science loss.

Preparations are complete for the start of the spring 2012 eclipse season; the season contains 21 eclipses and runs from Mar 16 to May 8. A real-time procedure was executed on Mar 15, in preparation for the start of the eclipse season, to dump and clear the EPS glitch counters. Chandra passed through the first eclipse of the season on Mar 16, with nominal power and thermal performance.

A real-time procedure was executed on Mar 15 to update the on-board ephemeris.

A Chandra image release was issued on Mar 14 describing observations of the galaxy cluster Abell 383. Two teams of researchers have used data from Chandra and other telescopes to map dark matter in 3-D in the galaxy cluster. Dark matter is invisible material that does not emit or absorb any type of light, but is detectable through its gravitational effects. Both teams found the dark matter to be distributed in the shape of a football, with the point close to the line of sight. For details see:

Of note this week, the Chandra Cycle 14 proposal deadline was Mar 15 at 6pm EDT.

The schedule of targets for the next week is shown below and includes observations of SN2011ht, which was accepted as a Target of Opportunity on Mar 5, of CH Cyg coordinated with HST, and of Sgr A* coordinated with several ground-based observatories.

Radiation Belts                     Mar 19
         Crab                   ACIS-S
         SDSSJ1034+2334         ACIS-S
         Comacluster            ACIS-I
         J1419.2+0756           ACIS-I       Mar 20
         SgrA*                  ACIS-S/HETG
         PSRJ1320-3512          ACIS-S
         Comacluster            ACIS-I
         Vega                   HRC-I        Mar 21
         Vega (2 obs)           HRC-S
         SgrA*                  ACIS-S/HETG
         Radiation Belts
         six-7                  ACIS-I
         Comacluster            ACIS-I
         SDSSJ1513+0855         ACIS-S       Mar 22
         SgrA*                  ACIS-S/HETG
         SDSSJ1457+0247         ACIS-S
         Comacluster            ACIS-I
         SDSSJ1532+3700         ACIS-S       Mar 23
         SgrA*                  ACIS-S/HETG
         SN2011ht               ACIS-S
         Radiation Belts                     Mar 24
         Z5768                  ACIS-I
         CHCyg                  ACIS-S
         J1214.0+2932           ACIS-I       Mar 25


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports